# A Diffusion-based Xray2MRI Model: Generating Pseudo-MRI Volumes From one Single X-ray

### v1

@@ Abstract
Knee osteoarthritis (KOA) is a prevalent musculoskeletal disorder, and X-rays are commonly used for its diagnosis due to their cost-effectiveness. Magnetic Resonance Imaging (MRI), on the other hand, offers detailed soft tissue visualization and has become a valuable supplementary diagnostic tool for KOA. Unfortunately, the high cost and limited accessibility of MRI hinder its widespread use, leaving many patients with KOA reliant solely on X-ray imaging. In this study, we introduce a novel diffusion-based Xray2MRI model capable of generating pseudo-MRI volumes from one single X-ray image. In addition to using X-rays as conditional input, our model integrates target depth, KOA probability distribution, and image intensity distribution modules to guide the synthesis process, ensuring that the generated corresponding slices accurately correspond to the anatomical structures. Experimental results demonstrate that by integrating information from X-rays with additional input data, our proposed approach is capable of generating pseudo-MRI sequences that approximate real MRI scans. Moreover, by increasing the inference times, the model achieves effective interpolation, further improving the continuity and smoothness of the generated MRI sequences, representing one promising initial attempt for cost-effective medical imaging solutions.
@@ 1 Introduction
Knee OsteoArthritis (KOA) is a progressive disease marked by the degeneration and damage of articular cartilage, changes at the joint margins, and reactive hyperplasia of the subchondral bone [1]. Various factors, such as age, obesity, stress, and trauma, can contribute to its development [2]. Those affected often endure severe pain and limited mobility, which can greatly diminish their quality of life and elevate the risk of chronic conditions like cardiovascular disease [3]. Despite extensive research, the precise cause of KOA remains unknown, and there is currently no cure [4]. Traditionally, X-ray imaging has been the most common method for diagnosing KOA due to its accessibility and ability to reveal changes in bone structure, such as joint space narrowing and osteophyte formation [5]. In 1957, Kellgren and Lawrence introduced the Kellgren-Lawrence (KL) grading system for assessing KOA [6] on X-ray. According to Table 1, this system categorizes KOA into five distinct grades, determined by the presence and severity of symptoms such as osteophytes and Joint Space Narrowing (JSN).
Table 1: Detailed Overview of the KL Grading System
Grade Severity Detailed description
KL-0 None Definitive absence of any osteoarthritis signs
KL-1 Doubtful Possible presence of initial osteophytic lipping
KL-2 Minimal Certain osteophytes formation and potential JSN
KL-3 Moderate Multiple moderate osteophytes, confirmed JSN, some
bone sclerosis, and potential bone end deformities
KL-4 Severe Large and numerous osteophytes, confirmed JSN, and
definitive deformation of bone ends
@@
However, X-ray imaging has limitations, particularly in visualizing soft tissue structures and early cartilage changes [7]. Magnetic Resonance Imaging (MRI), on the other hand, provides a more comprehensive assessment of the knee joint by offering detailed images of both bone and soft tissues, including cartilage, menisci, and synovium [8]. MRI can detect early signs of KOA that are not visible on X-rays, such as bone marrow lesions, synovitis, and early cartilage degeneration [9]. However, the widespread adoption of MRI in the clinical management of KOA is hindered by several challenges, most notably the high cost associated with MRI scans [10]. The expense of MRI technology, including the costs of the equipment, maintenance, and specialized personnel, makes it less accessible compared to traditional X-ray imaging [11]. This economic barrier is particularly significant in resource-limited settings and for patients without adequate health insurance coverage. In addition to the high financial cost, MRI scans require more time and patient cooperation, as they necessitate the patient to remain still for extended periods [12], which can be difficult for those experiencing severe pain or limited mobility due to KOA.
@@
To address these limitations, researchers have explored various techniques to enhance MRI reconstruction and reduce its dependency on traditional methods. Fessler et al. [13] provides an in-depth review of iterative algorithms used for improving MRI image reconstruction, particularly in scenarios where traditional inverse Fourier transforms are inadequate. The author highlights the limitations of conventional methods, especially when dealing with non-Cartesian sampling, under-sampled data, or the presence of nonlinear magnetic fields, and then focuses on model-based approaches, which utilize iterative methods that incorporate physical models of MRI acquisition, allowing for better image quality by addressing inhomogeneity and under-sampling issues. The authors demonstrate that model-based methods can significantly enhance image reconstruction, providing superior results compared to traditional techniques, though often at the cost of increased computation. In [14], Han et al. proposed leveraging the k-space to improve the efficiency and accuracy of MRI reconstruction. The authors introduced a framework inspired by the low-rank Hankel matrix completion approach, which directly interpolates missing k-space data and enhances the representation of MRI data by learning the underlying signal structure, resulting in high-quality images even from undersampled data. The proposed k-space deep learning approach significantly reduces the computational complexity and memory requirements. In [15], Lustig et al. leveraged compressed sensing to significantly accelerate MRI scan times without compromising image quality. The authors explore the inherent sparsity of MR images, both in the image domain and in transform domains such as spatial finite differences or wavelet coefficients. By randomly undersampling k-space and using nonlinear recovery schemes, the study shows that high-quality MR images can be reconstructed from incomplete data. This approach effectively reduces acquisition time and mitigates artefacts that arise from undersampling. The authors conclude that compressed sensing has the potential to substantially reduce scan times in clinical settings, making MRI more efficient and accessible.
@@
With the rapid advancements in deep learning technologies, the field of MRI reconstruction has seen significant growth [16] [17] [18]. Recently, combining multi-modal and diffusion-based models has become a new trend for image generation. In [19], Zhan et al. presented a novel medical generative model, MedM2G, designed to unify multiple medical modalities (e.g., CT, MRI, X-ray) and tasks (e.g., text-to-image, image-to-text, modality translation) into a single, efficient framework. The authors address the challenge of aligning and generating medical multi-modal data, particularly in the absence of large, well-paired datasets. They propose a central alignment strategy to efficiently align different modalities using text as the central modality and introduce medical visual invariant preservation to retain key clinical knowledge from each modality. The model leverages a multi-flow cross-guided diffusion process that facilitates adaptive interaction among modalities, enhancing the generation of accurate and high-quality medical images. In [20], Jiang et al. introduced a novel diffusion-based approach to multi-modal MRI synthesis. Traditional diffusion models for MRI synthesis typically operate in the original image domain, leading to high memory consumption and less effective handling of anatomical structures. CoLa-Diff addresses these challenges by operating in the latent space to reduce memory overhead while preserving anatomical details. The model introduces structural guidance using brain region masks and an auto-weight adaptation mechanism to balance multiple input modalities effectively during synthesis. The authors demonstrate that CoLa-Diff outperforms existing state-of-the-art models in multi-modal MRI synthesis tasks, showing superior results in terms of PSNR and SSIM. In [21], Kim et al. propose an advanced method for translating MRI images between different modalities. The authors introduce an Adaptive Latent Diffusion Model (ALDM) that leverages a novel multiple-switchable spatially adaptive normalization (MS-SPADE) block to transform source latent representations into target-like latent. This approach enables high-quality image synthesis across multiple modalities without requiring patch cropping, preserving the global information of the images. By conditioning the model with target modalities and applying the ALDM in a compressed latent space, the proposed method significantly reduces computational complexity while achieving superior results in both one-to-one and one-to-many image translation tasks. The authors highlight the model’s effectiveness in generating anatomically accurate and high-fidelity 3D medical images, outperforming existing methods in both qualitative and quantitative evaluations.
@@
    A novel CLDM-based framework, Xray2MRI, was proposed to generate the pseudo-MRI volumes from a single X-ray image.
@@
    In addition to the target depth, information about the KOA probability distribution and image intensity distribution from the X-ray is combined via an adaptive-weight module to guide the denoising process.
@@
    A new SSIM-based evaluation metric was specially developed. The multi-indicator assessment approach ensures both quantitative and qualitative validation of the synthesized MRI images for both image quality and clinical relevance.
@@
    The model showcases effective interpolation between slices by increasing the inference times (i.e., exceeding the depth of the original MRI volume), resulting in smoother transitions and improved image continuity.
@@ 2.1 Classical conditional latent diffusion model
Figure 2: The flowchart of the proposed approach, with black arrows indicating data flow, starts by extracting a slice at position $d$ from a real MRI sequence, denoted as $x_{d}$. This slice is encoded by the encoder $\mathcal{E}_{1}$ to produce a latent representation $z_{d}$, which is progressively noised following the distribution $q(z_{t}|z_{t-1})$, ultimately generating the noised latent representation $z_{T}$. The corresponding X-ray image $x_{c}$ is then used as a conditional input, encoded by $\mathcal{E}_{2}$, and concatenated with $z_{T}$ to form the initial input for the downsampling network. During the denoising process, the target depth $d$ is embedded at various stages of the downsampling phase. Concurrently, the KOA probability distribution $p$, obtained from a pre-trained KOA classifier $\mathcal{C}$, and the image intensity distribution $i$ from the X-ray image are integrated into the upsampling phase through an adaptive weighting module, providing dynamic guidance. This denoising process is repeated for $T-1$ iterations to progressively refine the latent representation, resulting in the denoised latent $\hat{z}_{d}$, which is decoded by $\mathcal{D}$ to reconstruct the MRI slice $\hat{x}_{d}$ at depth $d$. During inference, the process begins with $z_{T}$ and is repeated $c$ times, corresponding to the number of channels in the original MRI sequence. Finally, the full pseudo-MRI sequence $\hat{x}$ is obtained by stacking all the generated MRI slices in sequential order (i.e., $\{\hat{x}_{d=1}$, $\hat{x}_{d=2}$,…,$\hat{x}_{d=c}\}$).
@@ 2.1.1 Encode process
$z=\mathcal{E}(x)$ (1)
where $\mathcal{E}$ is an encoder which maps an image $x$ to a latent representation $z$.
@@ 2.1.3 Denoising and Sampling
$p_{\theta}(z_{0:T})=p(z_{T})\prod_{t=1}^{T}p_{\theta}(z_{t-1}\mid z_{t}),\quad p(z_{T})=\mathcal{N}(z_{T};0,\mathbf{I})$ (5)
where
$p_{\theta}(z_{t-1}\mid z_{t})=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t,y),\Sigma_{\theta}(z_{t},t,y))$ (6)
$p_{\theta}(z_{t-1}\mid z_{t})$ represents parameterized Gaussian distributions, the mean $\mu_{\theta}(z_{t},t,y)$ and variance $\Sigma_{\theta}(z_{t},t,y)$ with the condition $y$ are determined by trained learning networks.
@@ 2.1.4 Training process
$\mathcal{L}_{diff}=\mathbb{E}_{z\sim E(x),t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}(z_{t},t,c(x,y))||_{2}^{2}$ (8)
where $\mathbb{E}$ represents the expectation over the joint distribution of the original latent representation $z_{0}$ and Gaussian noise $\epsilon$.
@@ 2.2 Proposed approach
As presented in Fig. 2, our approach consists of three main modules: an auto-encoder module, a CLDM module, and a guidance module. $x_{d}$ is one slice of the original knee MRI sequence $x$ at the depth position $d$ with the size of $H\times W\times C$. $\mathcal{E}_{1}$ and $\mathcal{D}$ are encoder and decoder of the AutoencoderKL [25]. Unlike traditional autoencoders that map input data to a deterministic latent space, the AutoencoderKL introduces a probabilistic approach by encoding the input into a Gaussian distribution over the latent space. During training, the Kullback-Leibler (KL) divergence is used as a regularization term to ensure that the learned latent space closely follows a standard normal distribution, promoting smoother and more continuous transitions between latent representations. Specifically, the encoder operates at a resolution of 256$\times$256$\times$3 and begins with a convolutional layer configured with 128 channels and then proceeds through a series of stages, each with 2 residual blocks. The channel dimensions are multiplied according to the sequence 1, 2, 4, and 4. The decoder mirrors the structure of the encoder. The architecture is designed with no dropout, focusing on retaining all features without stochastic regularization during training. At the same time. the noised latent code $z_{T}$ is calculated following the Eq. 4:
$z_{T}=\sqrt{\bar{\alpha_{t}}}\mathcal{E}_{1}({x_{0}})+\sqrt{1-\bar{\alpha_{T}}}\epsilon$ (9)
@@
Next, the target depth $d$ was embedded into the downsampling stage to allow the model to condition the latent representation with spatial hierarchy early in the process. This ensures that depth-related features, such as different tissue layers and bone structures, are incorporated from the outset, preserving the anatomical coherence of the volume. In the upsampling stage, the KOA probability distribution $p$ and image intensity distribution $i$ were embedded to guide the synthesis of fine clinical features based on the spatial hierarchy established during downsampling, such as the presence of osteophytes and the size of the joint space. However, since the variation between slices in MRI is not uniform, different slices have varying levels of sensitivity to the guided information. Some slices may require more emphasis on specific anatomical features, while others may need more focus on KOA severity. To address this, we introduced an automatic weight adjustment mechanism that dynamically combines the information from the KOA probability distribution and the image intensity distribution, which ensures that the final MRI outputs maintain a higher level of accuracy and clinical relevance. The adaptive-weight module $w$ begins by independently adjusting the weights of the KOA probability distribution $p$ and the X-ray image intensity distribution $i$ using L2 norm, scaled by learnable parameters $\mu_{d}$ and $\mu_{p}$, respectively:
@@
Then, the calculated weights $G_{p}$ and $G_{i}$ are normalized as follows:
$\hat{G}_{p}=\frac{G_{p}}{\sqrt{G_{p}^{2}+\epsilon}},\quad\hat{G}_{i}=\frac{G_{i}}{\sqrt{G_{i}^{2}+\epsilon}}$ (12)
The KOA probability distribution $p$ and image intensity distribution $i$ are then adjusted by applying the normalized weights:
@@
As presented before, for the diffusion process, U-Net $\mathcal{U}$ is trained based on the Eq. 8:
@@
During the reverse process, the latent variable $z_{t-1}$ is sampled using the Denoising Diffusion Implicit Model (DDIM) [29], which enables deterministic or stochastic generation depending on the chosen schedule. DDIM refines the sampling process by approximating the reverse diffusion trajectory more efficiently than traditional diffusion models. Specifically, it leverages a non-Markovian process that skips some diffusion steps, allowing for faster inference without compromising the quality of the generated images. The sampling of $z_{t-1}$ is computed as follows:
@@
Finally, for each reconstructed slice at the depth $d$, denoted as $\hat{x}_{d}$, it is generated through a decoder $\mathcal{D}$. During the inference phase, this process is repeated $t$ times, corresponding to the channel (depth) $c$ of the original MRI volume. The final generated pseudo-MRI sequence $\hat{x}$ is formed by stacking all generated slices, represented as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{t=c}\}$. For more clarity, Algorithm 1 describes the entire process of the proposed approach.
Input: $x$, $c$, $t$, $x_{c}$, $\mathcal{E}_{2}$, $\mathcal{C}$ and all initial parameters
@@
             Get $\epsilon_{\theta},\hat{z}_{d}\leftarrow\mathcal{U}(z_{T},\mathcal{E}_{2}(x_{c}),t,w(p,i))$triangleright Eq. 16
@@
Input: $c$, $t$, $w$, $x_{c}$, $\mathcal{E}_{2}$, $\mathcal{C}$, $\mathcal{D}$ and $\mathcal{U}$
@@
       Get $p,i\leftarrow\mathcal{C}(x),intensity(x)$
       Get $\hat{z}_{d}\leftarrow\mathcal{U}(\epsilon,\mathcal{E}_{2}(x_{c}),t,w(p,i))$triangleright Eq. 16

### v2

@@ Abstract
Knee osteoarthritis (KOA) is a prevalent musculoskeletal disorder, and X-rays are commonly used for its diagnosis due to their cost-effectiveness. Magnetic Resonance Imaging (MRI), on the other hand, offers detailed soft tissue visualization and has become a valuable supplementary diagnostic tool for KOA. Unfortunately, the high cost and limited accessibility of MRI hinders its widespread use, leaving many patients with KOA to rely solely on X-ray imaging. In this study, we introduce a novel diffusion-based Xray2MRI model capable of generating pseudo-MRI volumes from a single X-ray image. In addition to using X-rays as conditional input, our model integrates target depth, KOA probability distribution, and image intensity distribution modules to guide the synthesis process, ensuring that the generated corresponding slices accurately correspond to the anatomical structures. Experimental results demonstrate that by integrating information from X-rays with additional input data, our proposed approach is capable of generating pseudo-MRI sequences that approximate real MRI scans. In addition, by increasing the number of inference steps, the model achieves effective interpolation, which further improves the continuity and smoothness of the generated MRI sequences, representing a promising first attempt at cost-effective medical imaging solutions. This study is available on https://zwang78.github.io/.
@@ 1 Introduction
Knee OsteoArthritis (KOA) is a progressive disease marked by the degeneration and damage of articular cartilage, changes at the joint margins, and reactive hyperplasia of the subchondral bone [1]. Various factors, such as age, obesity, stress, and trauma, can contribute to its development [2]. Those affected often endure severe pain and limited mobility, which can greatly diminish their quality of life and elevate the risk of chronic conditions like cardiovascular disease [3]. Despite extensive research, the precise cause of KOA remains unknown, and there is currently no cure [4].
@@
Traditionally, X-ray imaging has been the most common method for diagnosing KOA due to its accessibility and ability to reveal changes in bone structure, such as Joint Space Narrowing (JSN) and osteophyte formation [5]. To assess KOA on X-rays, five distinct grades (i.e., from G-0 to G-4), determined by the presence and severity of symptoms, are commonly used [6]. However, X-ray imaging has limitations, particularly in visualizing soft tissue structures and early cartilage changes. Magnetic Resonance Imaging (MRI), on the other hand, provides a more comprehensive assessment of the knee joint by offering detailed images of both bone and soft tissues, including cartilage, menisci, and synovium [7]. MRI can detect early signs of KOA that are not visible on X-rays, such as bone marrow lesions, synovitis, and early cartilage degeneration [8]. However, the widespread adoption of MRI in the clinical management of KOA is hindered by several challenges, most notably the high cost associated with MRI scans [9]. The expense of MRI technology, including the costs of the equipment, maintenance, and specialized personnel, makes it less accessible compared to traditional X-ray imaging [10]. This economic barrier is particularly significant in resource-limited settings and for patients without adequate health insurance coverage. In addition to the high financial cost, MRI scans require more time and patient cooperation, as they necessitate the patient to remain still for extended periods [11], which can be difficult for those experiencing severe pain or limited mobility due to KOA.
@@
To address these limitations, researchers have explored various techniques to enhance MRI reconstruction and reduce its dependency on traditional methods. Fessler et al. [12] provides an in-depth review of iterative algorithms used for improving MRI image reconstruction, particularly in scenarios where traditional inverse Fourier transforms are inadequate. The author highlights the limitations of conventional methods, especially when dealing with non-Cartesian sampling, under-sampled data, or the presence of nonlinear magnetic fields, and then focuses on model-based approaches, which utilize iterative methods that incorporate physical models of MRI acquisition, allowing for better image quality by addressing inhomogeneity and under-sampling issues. The authors demonstrate that model-based methods can significantly enhance image reconstruction, providing superior results compared to traditional techniques, though often at the cost of increased computation. In [13], Han et al. proposed leveraging the k-space to improve the efficiency and accuracy of MRI reconstruction. The authors introduced a framework inspired by the low-rank Hankel matrix completion approach, which directly interpolates missing k-space data and enhances the representation of MRI data by learning the underlying signal structure, resulting in high-quality images even from undersampled data. The proposed k-space deep learning approach significantly reduces the computational complexity and memory requirements. In [14], Lustig et al. leveraged compressed sensing to significantly accelerate MRI scan times without compromising image quality. The authors explore the inherent sparsity of MR images, both in the image domain and in transform domains. By randomly undersampling k-space and using nonlinear recovery schemes, the study shows that high-quality MR images can be reconstructed from incomplete data. This approach effectively reduces acquisition time and mitigates artefacts that arise from undersampling, which demonstrates that compressed sensing has the potential to substantially reduce scan times in clinical settings, making MRI more efficient and accessible.
@@
With the rapid advancements in deep learning technologies, the field of MRI reconstruction has seen significant growth [15] [16]. Recently, combining multi-modal and diffusion-based models has become a new trend for image generation. In [17], Zhan et al. presented a novel medical generative model, MedM2G, designed to unify multiple medical modalities (e.g., CT, MRI, X-ray) and tasks (e.g., text-to-image, image-to-text, modality translation) into a single, efficient framework. The authors address the challenge of aligning and generating medical multi-modal data, particularly in the absence of large, well-paired datasets. They propose a central alignment strategy to efficiently align different modalities using text as the central modality and introduce medical visual invariant preservation to retain key clinical knowledge from each modality. The model leverages a multi-flow cross-guided diffusion process that facilitates adaptive interaction among modalities, enhancing the generation of accurate and high-quality medical images. In [18], Jiang et al. introduced a novel diffusion-based approach to multi-modal MRI synthesis. Traditional diffusion models for MRI synthesis typically operate in the original image domain, leading to high memory consumption and less effective handling of anatomical structures. CoLa-Diff addresses these challenges by operating in the latent space to reduce memory overhead while preserving anatomical details. The model introduces structural guidance using brain region masks and an auto-weight adaptation mechanism to balance multiple input modalities effectively during synthesis. The authors demonstrate that CoLa-Diff outperforms existing state-of-the-art models in multi-modal MRI synthesis tasks, showing superior results in terms of Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM). In [19], Kim et al. propose an advanced method for translating MRI images between different modalities. The authors introduce an Adaptive Latent Diffusion Model (ALDM) that leverages a novel multiple-switchable spatially adaptive normalization (MS-SPADE) block to transform source latent representations into target-like latent. This approach enables high-quality image synthesis across multiple modalities without requiring patch cropping, preserving the global information of the images. By conditioning the model with target modalities and applying the ALDM in a compressed latent space, the proposed method significantly reduces computational complexity while achieving superior results in both one-to-one and one-to-many image translation tasks.
@@
    A novel CLDM-based framework, named Xray2MRI, is proposed to generate pseudo-MRI volumes from a single X-ray image.
@@
    In addition to the target depth, information about KOA probability distribution and image intensity distribution from the X-ray is combined via an adaptive-weight module to guide the denoising process.
@@
    A new SSIM-based evaluation metric is specially developed. The multi-indicator assessment approach ensures both quantitative and qualitative validation of the synthesized MRI images for both image quality and clinical relevance.
@@
    The model demonstrates effective interpolation between MRI slices by increasing the number of inference steps (i.e., exceeding the depth of the original MRI volume), resulting in smoother transitions and improved continuity between the generated slices.
@@ 2.1 Classical conditional latent diffusion model
Figure 2: The flowchart of the proposed approach, starts by extracting a slice at position $d$ from a real MRI sequence $x$, denoted as $x_{d}$. This slice is encoded by the encoder $\mathcal{E}_{1}$ to produce a latent representation $z_{0}$, which is progressively noised following the distribution $q(z_{t}|z_{t-1})$, ultimately generating the noised latent representation $z_{T}$. The corresponding X-ray image $x_{c}$ is then used as a conditional input, encoded by $\mathcal{E}_{2}$, and concatenated with $z_{T}$ to form the initial input for the downsampling network. During the denoising process, the target depth $d$ is embedded at various stages of the downsampling phase. Concurrently, the KOA probability distribution $p$, obtained from a pre-trained KOA classifier $\mathcal{C}$, and the image intensity distribution $i$ from the X-ray image are integrated into the upsampling phase through an adaptive weighting module, providing dynamic guidance. This denoising process is repeated for $T-1$ iterations to progressively refine the latent representation, resulting in the denoised latent $\hat{z}_{0}$, which is decoded by $\mathcal{D}$ to reconstruct the MRI slice $\hat{x}_{d}$ at depth $d$. During inference, the process begins with $z_{T}$ and is repeated $S$ times, corresponding to the number of channels $S$ in the original MRI sequence. Finally, the full pseudo-MRI sequence $\hat{x}$ is obtained by stacking all the generated MRI slices in sequential order (i.e., $\hat{x}$ = $\{\hat{x}_{d=1}$, $\hat{x}_{d=2}$,…,$\hat{x}_{d=S}\}$).
@@ 2.1.1 Encode process
$z_{0}=\mathcal{E}(x)$ (1)
where $\mathcal{E}$ is an encoder which maps an image $x$ to a latent representation $z_{0}$.
@@ 2.1.3 Denoising process
$p_{\theta}(z_{t-1}\mid z_{t},x_{c})=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t,x_{c}),\sigma_{\theta}^{2}(t))$ (5)
where $p_{\theta}(z_{t-1}\mid z_{t},x_{c})$ represents parameterized Gaussian distributions, the mean $\mu_{\theta}(z_{t},t,x_{c})$ and variance $\sigma_{\theta}^{2}(t)$ at each time step with the condition $x_{c}$ are determined by trained learning networks.
@@ 2.1.4 Training process
$\mathcal{L}_{diff}=\mathbb{E}_{z_{0}\sim\mathcal{E}(x),x_{c},t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}(z_{t},t,x_{c})||_{2}^{2}$ (7)
where $\mathbb{E}$ represents the expectation over all data points under the joint distribution of the conditioning information $x_{c}$, time steps $t$, and Gaussian noise $\epsilon$.
@@ 2.2 Proposed approach
As presented in Fig. 2, our approach consists of three main modules: an auto-encoder module, a CLDM module, and a guidance module. $x_{d}$ is a slice of the original knee MRI sequence $x$ at the depth position $d$ with the size of $H\times W\times S$. $\mathcal{E}_{1}$ and $\mathcal{D}$ are encoder and decoder of the AutoencoderKL [23]. Unlike traditional autoencoders that map input data to a deterministic latent space, the AutoencoderKL introduces a probabilistic approach by encoding the input into a Gaussian distribution over the latent space. During training, the Kullback-Leibler (KL) divergence is used as a regularization term to ensure that the learned latent space closely follows a standard normal distribution, promoting smoother and more continuous transitions between latent representations. Specifically, the encoder operates at a resolution of 256$\times$256$\times$3 and begins with a convolutional layer configured with 128 channels and then proceeds through a series of stages, each with 2 residual blocks. The channel dimensions are multiplied according to the sequence 1, 2, 4, and 4. The decoder mirrors the structure of the encoder. The architecture is designed with no dropout, focusing on retaining all features without stochastic regularization during training. Simultaneously, the noised latent code $z_{T}$ is calculated following Eq. 4:
$z_{T}=\sqrt{\bar{\alpha_{t}}}\mathcal{E}_{1}({x_{d}})+\sqrt{1-\bar{\alpha}_{T}}\epsilon$ (8)
@@
Next, the target depth $d$ was embedded into the downsampling stage to allow the model to condition the latent representation with spatial hierarchy early in the process. This ensures that depth-related features, such as different tissue layers and bone structures, are incorporated from the outset, preserving the anatomical coherence of the volume. In the upsampling stage, the KOA probability distribution $p$ and image intensity distribution $i$ were embedded to guide the synthesis of fine clinical features based on the spatial hierarchy established during downsampling, such as the presence of osteophytes and the size of the joint space. However, since the variation between slices in MRI is not uniform, different slices have varying levels of sensitivity to the guided information. Some slices may require more emphasis on specific anatomical features, while others may need more focus on KOA severity. To address this, we introduce an automatic weight adjustment mechanism that dynamically combines the information from the KOA probability distribution $p$ and the image intensity distribution $i$ to ensure that the final MRI outputs maintain a higher level of accuracy and clinical relevance by properly leveraging the complementary strengths of both distributions. The adaptive-weight module $w$ begins by independently adjusting the weights of $p$ and $i$ using the $L^{2}$-norm, scaled by the learnable parameters $\mu_{p}$ and $\mu_{i}$, respectively:
@@
Next, the calculated weights $G_{p}$ and $G_{i}$ are normalized as follows:
$\hat{G}_{p}=\frac{G_{p}}{G_{p}+G_{i}+\epsilon},\quad\hat{G}_{i}=\frac{G_{i}}{G_{p}+G_{i}+\epsilon}$ (11)
where $\epsilon$ is a small positive constant (e.g., 1e-08) added for numerical stability, preventing division by zero.
Then, $p$ and $i$ are adjusted by applying the normalized weights:
@@
As presented before, the diffusion model is implemented through a U-Net $\mathcal{U}$, which is trained based on Eq. 7. In this study, the goal function is computed as follows:
@@
During the denoising process, the latent variable $z_{t-1}$ is sampled using the Denoising Diffusion Implicit Model (DDIM) [27], which enables deterministic or stochastic generation depending on the chosen schedule. DDIM refines the sampling process by approximating the reverse diffusion trajectory more efficiently than traditional diffusion models. Specifically, it leverages a non-Markovian process that skips some diffusion steps, allowing for faster inference without compromising the quality of the generated images. The sampling of $z_{t-1}$ is computed as follows:
@@
Finally, for each reconstructed slice at the depth $d$, denoted as $\hat{x}_{d}$, it is generated through a decoder $\mathcal{D}$. During the inference phase, this process is repeated $S$ times, corresponding to the channel (depth) $S$ of the original MRI volume. The final generated pseudo-MRI sequence $\hat{x}$ is formed by stacking all generated slices, represented as $\hat{x}=\{\hat{x}_{0},\hat{x}_{1},\dots,\hat{x}_{s-1}\}$. For more clarity, Algorithm 1 and Algorithm 2 describe the training and inference process of the proposed approach, respectively.
Input: $x$, $t$, $S$, $x_{c}$, $\mathcal{E}_{2}$, $\mathcal{C}$ and all initial parameters
@@
             Get $p,i\leftarrow\mathcal{C}(x_{c}),intensity(x_{c})$
             Get $\epsilon_{\theta},\hat{z}_{d}\leftarrow\mathcal{U}(z_{T},\mathcal{E}_{2}(x_{c}),t,w(p,i))$
@@
Input: $t$, $w$, $S$, $x_{c}$, $\mathcal{E}_{2}$, $\mathcal{C}$, $\mathcal{D}$ and $\mathcal{U}$
@@
for $d$ in $\{0,1,...,S-1\}$ do
       Get $p,i\leftarrow\mathcal{C}(x_{c}),intensity(x_{c})$
       Get $\hat{z}_{d}\leftarrow\mathcal{U}(\epsilon,\mathcal{E}_{2}(x_{c}),t,w(p,i))$